# Birch SGD: A Tree Graph Framework for Local and Asynchronous SGD Methods

### v2

@@ Appendix B Notations
:=$\mathbb{N}:=\{1,2,\dots\};$ $\left\|x\right\|$ is the output of the standard Euclidean norm for all $x\in\mathbb{R}^{d}$; $\left\langle x,y\right\rangle=\sum_{i=1}^{d}x_{i}y_{i}$ is the standard dot product; $g=\mathcal{O}(f):$ exist $C>0$ such that $g(z)\leq C\times f(z)$ for all $z\in\mathcal{Z};$ $g=\Omega(f):$ exist $C>0$ such that $g(z)\geq C\times f(z)$ for all $z\in\mathcal{Z};$ $g=\Theta(f):$ $g=\mathcal{O}(f)$ and $g=\Omega(f);$ $g=\widetilde{\Theta}(f):$ the same as $g=\Omega(f)$ but up to logarithmic factors; $a\vee b:=\max\{a,b\}.$
@@ C.1 Setup
The experiments were prepared in Python. The distributed environment was simulated with the Simpy Python library (Matloff,, 2008). There are two hardware setups:
@@ C.2 Experiments with logistic regression
We begin our experiments with one the simplest ML problems—logistic regression on the MNIST dataset LeCun et al., (2010). In this setting, we evaluate three different numbers of workers: $n\in\{16,64,256\}$. We use the standard linear model with the logistic loss.
@@ C.3 Experiments with ResNet18 and image classification
We test the algorithms on the CIFAR10 (Krizhevsky et al.,, 2009) image recognition task with the ResNet18 (He et al.,, 2016) deep neural network. For ResNet18, we similarly report the best convergence results from a grid search over the parameters listed in Table 6, using a setup with $n=8$ workers. Results for all algorithms across the four regimes are presented in Figure 11.
@@ C.4 Experiments with GPT2 and token prediction
We also evaluate the algorithms on the Wikitext-2 (Merity et al.,, 2016) next token prediction task with GPT2 (Radford et al.,, 2019). For GPT2, we evaluate all four regimes using a setup with $n=8$ workers. To achieve faster and more robust convergence, we use the AdamW normalization strategy [^6] Instead of the SGD step $w^{k+1}=w^{k}-\gamma g^{k},$ where $g^{k}$ is a descent direction, we use the AdamW strategy with $g^{k}$. only in these experiments with GPT2. The resulting convergence curves are shown in Figure 12. Once again, the results are similar to those of the previous experiments. Due to hardware limitations, a narrower grid search range is used; therefore, it is possible that convergence could be further improved with a more extensive search.
@@
C.5 Parameters of the experiments
@@
Synchronized SGD $\gamma$ range: $[2^{-5},2^{4}]$
Rennala SGD $\gamma$ range: $[2^{-15},2^{-3}]$
@@
Local SGD $\gamma$ range: $[2^{-15},2^{-3}]$
@@
Ringmaster ASGD $\gamma$ range: $[2^{-15},2^{1}]$
@@
Async-Local SGD $\gamma$ range: $[2^{-10},2^{1}]$
@@
Synchronized SGD $\gamma$ range: $[2^{-5},2^{4}]$
Rennala SGD $\gamma$ range: $[2^{-15},2^{-3}]$
@@
Local SGD $\gamma$ range: $[2^{-15},2^{-3}]$
@@
Ringmaster ASGD $\gamma$ range: $[2^{-15},2^{1}]$
@@
Async-Local SGD $\gamma$ range: $[2^{-10},2^{1}]$
@@
Synchronized SGD $\gamma$ range: $[2^{-5},2^{4}]$
Rennala SGD $\gamma$ range: $[2^{-15},2^{-3}]$
@@
Local SGD $\gamma$ range: $[2^{-15},2^{-3}]$
@@
Ringmaster ASGD $\gamma$ range: $[2^{-15},2^{1}]$
@@
Async-Local SGD $\gamma$ range: $[2^{-10},2^{1}]$
@@
Synchronized SGD $\gamma$ range: $[2^{-6},2^{-3}]$
Rennala SGD $\gamma$ range: $[2^{-6},2^{-3}]$
@@
Local SGD $\gamma$ range: $[2^{-6},2^{-3}]$
@@
Ringmaster ASGD $\gamma$ range: $[2^{-6},2^{-3}]$
@@
Async-Local SGD $\gamma$ range: $[2^{-6},2^{-3}]$
@@
Synchronized SGD $\gamma$ range: $[2^{-11},2^{-10}]$
Rennala SGD $\gamma$ range: $[2^{-11},2^{-10}]$
@@
Local SGD $\gamma$ range: $[2^{-11},2^{-10}]$
@@
Ringmaster ASGD $\gamma$ range: $[2^{-11},2^{-10}]$
@@
Async-Local SGD $\gamma$ range: $[2^{-11},2^{-10}]$
@@
$x^{k}-z^{k}=-\gamma\left(\sum_{i=p}^{k-1}\nabla f(z^{i};\xi^{i})-\sum_{(w,\xi)\in S^{k}}\nabla f(w;\xi)\right)=-\gamma\sum_{j\in\bar{S}^{k}}\nabla f(z^{i};\xi^{i}),$ (9)
@@
[…]
@@
${\mathbb{E}}\left[f(x^{K-1})-f^{*}\right]$ $\leq f(x^{0})-f^{*}-\frac{\gamma}{2}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\left\|\nabla f(x^{k})\right\|^{2}\right]-\frac{\gamma}{4}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\left\|\nabla f(z^{k})\right\|^{2}\right]+KL\gamma^{2}\sigma^{2}$
@@
because $\gamma\leq\frac{1}{2LR}.$ Finally, since ${\mathbb{E}}\left[f(x^{K-1})-f^{*}\right]\geq 0,$
@@ E.1 Vanilla SGD
While the iteration rate of Vanilla SGD is well-known (Lan,, 2020), we prove its convergence using our new framework for clarity.
@@ E.2 Rennala SGD
We now apply Theorem 2.4 to Rennala SGD. The iteration rate of Rennala SGD is also well-known (Tyurin and Richtárik,, 2023), but we provide a proof for completness. Rennala SGD can be formally described as follows:
Algorithm 4 Rennala SGD (Tyurin and Richtárik,, 2023)
@@
$x^{1}$ $=x^{0}-\gamma\nabla f(x^{0};\xi^{0}),\quad$ $\dots$,$\quad$ , $x^{B}=x^{B-1}-\gamma\nabla f(x^{0};\xi^{B-1}),$ (12)
$x^{B+1}$ $=x^{B}-\gamma\nabla f(x^{B};\xi^{B}),\quad$ $\dots$,$\quad$ , $x^{2B}=x^{2B-1}-\gamma\nabla f(x^{B};\xi^{2B-1}),\quad\dots,$
@@
The maximum tree distance between $x^{k}$ and $z^{k}$ is $B-1$. Thus, $R=B-1$ in Condition 3. ∎
@@ E.3 Local SGD
[…]
@@
[…]
@@
Although the proof may seem technical due to the heavy notation in (14), it is actually straightforward when you refer to Figure 4. This figure clearly shows that all conditions of Theorem 2.4 are satisfied with $R=B-1$ because $\sum_{i=1}^{n}M_{i}=B$ in every global iteration. The condition $\sum_{i=1}^{n}M_{i}=B$ helps us to insure that the maximum tree distance $\sup_{k\geq 0}\textnormal{dist}(x^{k},z^{k})\leq B-1.$
@@
because the maximum number of edges to the common closest ancestor can not exit $B-1$. ∎

### v3

@@ Appendix B Notations
$\mathbb{N}:=\{1,2,\dots\};$ $\left\|x\right\|$ is the output of the standard Euclidean norm for all $x\in\mathbb{R}^{d}$; $\left\langle x,y\right\rangle=\sum_{i=1}^{d}x_{i}y_{i}$ is the standard dot product; $g=\mathcal{O}(f):$ exist $C>0$ such that $g(z)\leq C\times f(z)$ for all $z\in\mathcal{Z};$ $g=\Omega(f):$ exist $C>0$ such that $g(z)\geq C\times f(z)$ for all $z\in\mathcal{Z};$ $g=\Theta(f):$ $g=\mathcal{O}(f)$ and $g=\Omega(f);$ $g=\widetilde{\Theta}(f):$ the same as $g=\Omega(f)$ but up to logarithmic factors; $a\vee b:=\max\{a,b\}.$
@@ C.1 Setup
The experiments were prepared in Python. The distributed environment was simulated with the Simpy Python library (Matloff, 2008). There are two hardware setups:
@@ C.2 Experiments with logistic regression
We begin our experiments with one the simplest ML problems—logistic regression on the MNIST dataset LeCun et al. (2010). In this setting, we evaluate three different numbers of workers: $n\in\{16,64,256\}$. We use the standard linear model with the logistic loss.
@@ C.3 Experiments with ResNet18 and image classification
We test the algorithms on the CIFAR10 (Krizhevsky et al., 2009) image recognition task with the ResNet18 (He et al., 2016) deep neural network. For ResNet18, we similarly report the best convergence results from a grid search over the parameters listed in Table 6, using a setup with $n=8$ workers. Results for all algorithms across the four regimes are presented in Figure 11.
@@ C.4 Experiments with GPT2 and token prediction
We also evaluate the algorithms on the Wikitext-2 (Merity et al., 2016) next token prediction task with GPT2 (Radford et al., 2019). For GPT2, we evaluate all four regimes using a setup with $n=8$ workers. To achieve faster and more robust convergence, we use the AdamW normalization strategy [^6] Instead of the SGD step $w^{k+1}=w^{k}-\gamma g^{k},$ where $g^{k}$ is a descent direction, we use the AdamW strategy with $g^{k}$. only in these experiments with GPT2. The resulting convergence curves are shown in Figure 12. Once again, the results are similar to those of the previous experiments. Due to hardware limitations, a narrower grid search range is used; therefore, it is possible that convergence could be further improved with a more extensive search.
@@
C.5 Experiments with Cycle SGD and peak bandwidth
Figure 13: Comparison of methods in the setup, where the communication time depends on the number of synchronized workers. It takes $h_{i}=10$ seconds to compute a stochastic gradient, and the communication time is $1.5\times k$ seconds, where $k$ is the number of synchronized workers.
In Figure 13, we present the comparison of Cycle SGD with other methods to verify the advantageous theoretical property of Cycle SGD. For this particular regime, when the computational times are the same, all methods except Cycle SGD require all the $64$ workers to synchronize at the same time, leading to high peak bandwidth. At the same time, Cycle SGD synchronizes only $8$ workers and reduces the communication time and the total time complexity, which we observe in Figure 13 with logistic regression on the MNIST dataset.
C.6 Sensitivity to the choice of $B$ in Rennala SGD
Figure 14: Slow Communications regime: $h_{i}=10$, $\tau_{i}=100$
In this section, we aim to understand the sensitivity of the choice of the value $B$ in Rennala SGD. From the proof of Theorem E.2, we know that $R:=\sup_{k\geq 0}\textnormal{dist}(x^{k},z^{k})=\Theta(B).$ By increasing $B=\Theta(R),$ we can examine whether the method starts to slow down. This slowdown is expected due to the discussion in Section 4. In Figure 14, we present our results on logistic regression problem on the MNIST dataset and $n=100$ workers, where we see that if we take $B$ too large, the method slows down accordingly.
C.7 Parameters of the experiments
@@
Synchronized SGD $\gamma$ range: $2^{-5},2^{4}]$
Rennala SGD $\gamma$ range: $2^{-15},2^{-3}]$
@@
Local SGD $\gamma$ range: $2^{-15},2^{-3}]$
@@
Ringmaster ASGD $\gamma$ range: $2^{-15},2^{1}]$
@@
Async-Local SGD $\gamma$ range: $2^{-10},2^{1}]$
@@
Synchronized SGD $\gamma$ range: $2^{-5},2^{4}]$
Rennala SGD $\gamma$ range: $2^{-15},2^{-3}]$
@@
Local SGD $\gamma$ range: $2^{-15},2^{-3}]$
@@
Ringmaster ASGD $\gamma$ range: $2^{-15},2^{1}]$
@@
Async-Local SGD $\gamma$ range: $2^{-10},2^{1}]$
@@
Synchronized SGD $\gamma$ range: $2^{-5},2^{4}]$
Rennala SGD $\gamma$ range: $2^{-15},2^{-3}]$
@@
Local SGD $\gamma$ range: $2^{-15},2^{-3}]$
@@
Ringmaster ASGD $\gamma$ range: $2^{-15},2^{1}]$
@@
Async-Local SGD $\gamma$ range: $2^{-10},2^{1}]$
@@
Synchronized SGD $\gamma$ range: $2^{-6},2^{-3}]$
Rennala SGD $\gamma$ range: $2^{-6},2^{-3}]$
@@
Local SGD $\gamma$ range: $2^{-6},2^{-3}]$
@@
Ringmaster ASGD $\gamma$ range: $2^{-6},2^{-3}]$
@@
Async-Local SGD $\gamma$ range: $2^{-6},2^{-3}]$
@@
Synchronized SGD $\gamma$ range: $2^{-11},2^{-10}]$
Rennala SGD $\gamma$ range: $2^{-11},2^{-10}]$
@@
Local SGD $\gamma$ range: $2^{-11},2^{-10}]$
@@
Ringmaster ASGD $\gamma$ range: $2^{-11},2^{-10}]$
@@
Async-Local SGD $\gamma$ range: $2^{-11},2^{-10}]$
@@
D.1 Proof technique and reasons for choosing the conditions
Before we state the main theorem and provide the proof, let us explain the intuition, the novelty, and how we identified the conditions of the theorem. The proof of the result is given in Section D.2 and is relatively compact. We believe that the simplicity of our result, together with its ability to unify methods, constitutes an important contribution to the optimization community. While the initial part of the proof follows the same structure as in most related works, starting from (7), our treatment of the staleness term $\|x^{k}-z^{k}\|,$ which naturally arises from the step $x^{k+1}=x^{k}-\gamma\nabla f(z^{k};\xi^{k}),$ is novel.
After many attempts to develop a universal theory, let us illustrate how we arrived at our conditions. Looking at Figure 3, which provides all possible relations between $x^{k}$ and $z^{k},$ one can easily get
$\left\|x^{k}-z^{k}\right\|=\gamma\left\|{\color[rgb]{0.09765625,0.09765625,0.75390625}\definecolor[named]{pgfstrokecolor}{rgb}{0.09765625,0.09765625,0.75390625}\sum_{i=p}^{k-1}\nabla f(z^{i};\xi^{i})}-{\color[rgb]{0.75390625,0.09765625,0.09765625}\definecolor[named]{pgfstrokecolor}{rgb}{0.75390625,0.09765625,0.09765625}\sum_{(w,\xi)\in S^{k}}\nabla f(w;\xi)}\right\|.$
First, we noticed that any reasonable method should utilize $\color[rgb]{0.75390625,0.09765625,0.09765625}\definecolor[named]{pgfstrokecolor}{rgb}{0.75390625,0.09765625,0.09765625}\sum_{(w,\xi)\in S^{k}}\nabla f(w;\xi)$} in the computation of $z^{k}$ before applying $\nabla f(z^{k};\xi^{k})$ (see the previous discussion about Condition 2 in Section 2.1). This implies ${\color[rgb]{0.75390625,0.09765625,0.09765625}\definecolor[named]{pgfstrokecolor}{rgb}{0.75390625,0.09765625,0.09765625}\{(w;\xi)\}_{(w,\xi)\in S^{k}}}\subseteq{\color[rgb]{0.09765625,0.09765625,0.75390625}\definecolor[named]{pgfstrokecolor}{rgb}{0.09765625,0.09765625,0.75390625}\{(z^{i};\xi^{i})\}_{i=p}^{k-1}},$ leading to the following identity:
$\left\|x^{k}-z^{k}\right\|=\gamma\left\|\sum_{j\in\bar{S}^{k}}\nabla f(z^{j};\xi^{j})\right\|$
for some set $\bar{S}^{k}\subseteq{\color[rgb]{0.09765625,0.09765625,0.75390625}\definecolor[named]{pgfstrokecolor}{rgb}{0.09765625,0.09765625,0.75390625}\{p,\dots,k-1\}}$ such that $\bar{S}^{k}\cup S^{k}={\color[rgb]{0.09765625,0.09765625,0.75390625}\definecolor[named]{pgfstrokecolor}{rgb}{0.09765625,0.09765625,0.75390625}\{p,\dots,k-1\}}.$ The identity says that the distance is roughly proportional to the number $\left|\bar{S}^{k}\right|$ of stochastic gradients applied after $x^{p}$ and before $z^{k},$ which is tightly bounded by the tree distance from $x^{k}$ to the common ancestor $x^{p},$ i.e., it is bounded by $\left|{\color[rgb]{0.09765625,0.09765625,0.75390625}\definecolor[named]{pgfstrokecolor}{rgb}{0.09765625,0.09765625,0.75390625}\{p,\dots,k-1\}}\right|$ since $\bar{S}^{k}\subseteq{\color[rgb]{0.09765625,0.09765625,0.75390625}\definecolor[named]{pgfstrokecolor}{rgb}{0.09765625,0.09765625,0.75390625}\{p,\dots,k-1\}}.$
Under Condition 2, notice that $\left|{\color[rgb]{0.09765625,0.09765625,0.75390625}\definecolor[named]{pgfstrokecolor}{rgb}{0.09765625,0.09765625,0.75390625}\{p,\dots,k-1\}}\right|=\max\{\left|{\color[rgb]{0.09765625,0.09765625,0.75390625}\definecolor[named]{pgfstrokecolor}{rgb}{0.09765625,0.09765625,0.75390625}\{p,\dots,k-1\}}\right|,\left|S^{k}\right|\}=:\textnormal{dist}(x^{k},z^{k}),$ where we use $S^{k}\subseteq{\color[rgb]{0.09765625,0.09765625,0.75390625}\definecolor[named]{pgfstrokecolor}{rgb}{0.09765625,0.09765625,0.75390625}\{p,\dots,k-1\}}$ and Definition 2.2. Thus, to get a bound for $\|x^{k}-z^{k}\|,$ it is natural to introduce Condition 3, which allows us to conclude that $\left|\bar{S}^{k}\right|\leq\textnormal{dist}(x^{k},z^{k})\leq R.$ It remains to use classical mathematical tools to obtain
${\mathbb{E}}\left[\left\|x^{k}-z^{k}\right\|^{2}\right]\leq 2\gamma^{2}R\sum_{j=k-R}^{k-1}{\mathbb{E}}\left[\left\|\nabla f(z^{j})\right\|^{2}\right]+2\gamma^{2}R\sigma^{2},$
where the first term will be canceled by the corresponding term $\frac{\gamma}{4}{\mathbb{E}}\left[\left\|\nabla f(z^{k})\right\|^{2}\right]$ from (6).
D.2 Full proof
@@
$x^{k}-z^{k}=-\gamma\left(\sum_{i=p}^{k-1}\nabla f(z^{i};\xi^{i})-\sum_{(w,\xi)\in S^{k}}\nabla f(w;\xi)\right)=-\gamma\sum_{j\in\bar{S}^{k}}\nabla f(z^{j};\xi^{j}),$ (9)
@@
$\overset{\textnormal{Jensen's ineq.}}{\leq}2L^{2}\gamma^{2}\left|\bar{S}^{k}\right|\sum_{j\in\bar{S}^{k}}{\mathbb{E}}\left[\left\|\nabla f(z^{j})\right\|^{2}\right]+2L^{2}\gamma^{2}\left|\bar{S}^{k}\right|\sigma^{2}.$
@@
${\mathbb{E}}\left[f(x^{K})-f^{*}\right]$ $\leq f(x^{0})-f^{*}-\frac{\gamma}{2}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\left\|\nabla f(x^{k})\right\|^{2}\right]-\frac{\gamma}{4}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\left\|\nabla f(z^{k})\right\|^{2}\right]+KL\gamma^{2}\sigma^{2}$
@@
because $\gamma\leq\frac{1}{2LR}.$ Finally, since ${\mathbb{E}}\left[f(x^{K})-f^{*}\right]\geq 0,$
@@ E.1 Vanilla SGD
While the iteration rate of Vanilla SGD is well-known (Lan, 2020), we prove its convergence using our new framework for clarity.
@@ E.2 Rennala SGD
We now apply Theorem 2.4 to Rennala SGD. The iteration rate of Rennala SGD is also well-known (Tyurin and Richtárik, 2023), but we provide a proof for completness. Rennala SGD can be formally described as follows:
Algorithm 4 Rennala SGD (Tyurin and Richtárik, 2023)
@@
$x^{1}$ $=x^{0}-\gamma\nabla f(x^{0};\xi^{0}),\quad$ $\dots,\quad$ $x^{B}=x^{B-1}-\gamma\nabla f(x^{0};\xi^{B-1}),$ (12)
$x^{B+1}$ $=x^{B}-\gamma\nabla f(x^{B};\xi^{B}),\quad$ $\dots,\quad$ $x^{2B}=x^{2B-1}-\gamma\nabla f(x^{B};\xi^{2B-1}),\quad\dots,$
@@
The maximum tree distance between $x^{k}$ and $z^{k}$ is B1B-1. Thus, R=B1R=B-1 in Condition 3. ∎
@@ E.3 Local SGD
$\begin{aligned} &z^{k,1}_{i}=z^{k,0}_{i}-\gamma\nabla f(x^{N_{k}};\eta^{k,0}_{i}),\\ &z^{k,2}_{i}=z^{k,1}_{i}-\gamma\nabla f(z^{k,1}_{i};\eta^{k,1}_{i}),\\ &\qquad\vdots\\ &z^{k,M_{i}}_{i}=z^{k,M_{i}-1}_{i}-\gamma\nabla f(z^{k,M_{i}-1}_{i};\eta^{k,M_{i}-1}_{i}),\end{aligned}$ (13)
@@
$\begin{aligned} x^{N_{k}+1}&=x^{N_{k}}-\gamma\nabla f(z^{k,0}_{1};\eta^{k,0}_{1}),\\ &\vdots\\ x^{N_{k}+M_{1}}&=x^{N_{k}+M_{1}-1}-\gamma\nabla f(z^{k,M_{1}-1}_{1};\eta^{k,M_{1}-1}_{1}),\\ x^{N_{k}+M_{1}+1}&=x^{N_{k}+M_{1}}-\gamma\nabla f(z^{k,0}_{2};\eta^{k,0}_{2}),\\ &\vdots\\ x^{N_{k}+M_{1}+M_{2}}&=x^{N_{k}+M_{1}+M_{2}-1}-\gamma\nabla f(z^{k,M_{2}-1}_{2};\eta^{k,M_{2}-1}_{2}),\\ &\vdots\\ x^{N_{k+1}}&=x^{N_{k}+\sum_{i=1}^{n}M_{i}-1}-\gamma\nabla f(z^{k,M_{n}-1}_{n};\eta^{k,M_{n}-1}_{n}).\end{aligned}$ (14)
@@
Although the proof may seem technical due to the heavy notation in (14), it is actually straightforward when you refer to Figure 4. This figure clearly shows that all conditions of Theorem 2.4 are satisfied with R=B1R=B-1 because $\sum_{i=1}^{n}M_{i}=B$ in every global iteration. The condition $\sum_{i=1}^{n}M_{i}=B$ helps us to insure that the maximum tree distance $\sup_{k\geq 0}\textnormal{dist}(x^{k},z^{k})\leq B-1.$
@@
because the maximum number of edges to the common closest ancestor can not exit B1B-1. ∎